# reference.bib

### v1

@@ Abstract
Optimization for robot control tasks, spanning various methodologies, includes Model Predictive Control (MPC). However, the complexity of the system, such as non-convex and non-differentiable cost functions and prolonged planning horizons often drastically increases the computation time, limiting MPC’s real-world applicability. Prior works in speeding up the optimization have limitations on solving convex problem and generalizing to hold out domains. To overcome this challenge, we develop a novel framework aiming at expediting optimization processes. In our framework, we combine offline self-supervised learning and online fine-tuning through reinforcement learning to improve the control performance and reduce optimization time.We demonstrate the effectiveness of our method on a novel, challenging Formula-1-track driving task, achieving $3.9\%$ higher performance in optimization time and $3.6\%$ higher performance in tracking accuracy on challenging holdout tracks.
@@ I INTRODUCTION
Iterative control optimization algorithms have been widely adopted to control the dynamic systems such as autonomous vehicles [borrelli2005mpc, karnchanachari2020practical, kong2015kinematic, kim2022smooth], aircraft [bauersfeld2021mpc, jadbabaie2002control], humanoid robots [kuindersma2016optimization], etc. These optimization paradigms often entail dealing with complex systems featuring constraints, high dimensional solution space, and sudden state changes. Due to the inherent complexity of such systems, finding a good-enough solution (within some tolerance) in a single attempt can be exceedingly challenging for the optimizer. Consequently, solvers adopt an iterative approach to gradually converge towards a satisfactory solution. Typically, these algorithms start with an initial guess of the control inputs (e.g., joint accelerations) and employs solvers to iteratively minimize the cost function [sacks2023learning]. However, the optimization process can be significantly complicated by factors, such as the non-convex and non-differentiable cost function [rawlings1994nonlinear, eren2017model]. Consequently, the optimization time of MPC is typically a bottleneck in its real-world applications [bouzidi2023learning, lembono2020learning, mansard2018using, richter2009real]. Addressing this challenge is paramount for enhancing the practical utility of MPC in various domains.
@@
Figure 1: Overview of our proposed algorithm. The first two blocks denote the two-phase training framework. In the first phase, we collect expert MPC demonstrations and train a warm-start policy using behavior cloning to speed up MPC. In the second phase, we fine-tune this policy within an online training framework to enhance its performance and generalizability. During testing, the proposed framework is evaluated on both training tracks and challenging holdout tracks, as demonstrated in the third block.
@@
Given the iterative nature of the solvers used in optimization for robot control tasks, providing the solver with a better initial guess can expedite the optimization process. This process, known as warm starting, involves initializing the optimization algorithm with a solution that is closer to the optimal solution than a randomly chosen starting point. The intuition behind warm starting is that, at each iteration, the solver refines its solution based on previous iterations. As shown in Figure 2, a closer initial guess reduces the distance the optimization algorithm needs to traverse to converge to the optimal solution. By starting closer to the optimum, the solver can often avoid lengthy exploration of the solution space, leading to faster convergence and reduced computational effort. In this paper, we choose to test our proposed algorithm to warm start Model Predictive Control (MPC), which optimizes a trajectory over a predefined horizon.
@@
To expedite the optimization process of MPC, various approaches have been adopted. However, they have fallen short on holdout domains, or have limitations in dealing with sudden state changes. Traditionally, one common technique involves utilizing the MPC solution from the previous sampling instance as the initial guess for the current control step [zeilinger2011real, pan2020imitation]. However, this approach falls short when faced with sudden state changes (e.g., the vehicle approaches a sharp turn). Another method involves maintaining a memory buffer to store historical MPC solutions, from which a suitable initial guess can be retrieved for future planning steps [mansard2018using, marcucci2020warm]. However, this approach may lack zero-shot generalizability. As the system’s state becomes heterogeneous, the size of the memory buffer needs to increase to maintain effectiveness. Moreover, searching for the initial guess within the memory buffer scales at least linearly with its size, potentially causing time-consuming operations. Alternatively, a learning-based methods have been proposed by Klaučo et al. to utilize k-NN classifier to classify the solution space into different active sets from which the solver searches for a solution [klauvco2019machine]. However, this method falls short in providing a precise warm-started initial guess and is not capable of dealing with challenging control tasks featuring long planning horizon and heterogeneous observations. Additionally, this approach is limited to solving strictly convex quadratic programs, imposing significant constraints on its applicability and generalizability given that many real-world control problems lack a linear or differentiable objective function and dynamics model. Thus, there is a pressing need for innovative approaches that can effectively address the challenges posed by the inherent complexity and non-linearity of real-world control systems.
@@
In this paper, we propose a two-phase learning framework to learn a warm-start policy that provides a better initial guess for iterative control optimization algorithms (such as MPC) to reduce the optimization time. Our proposed algorithm is depicted in Figure 1, which includes a two-phase training framework and an additional holdout testing phase. An advantage of utilizing a learned policy to initialize the MPC, as opposed to learning a policy for end-to-end control of the system, is that it maintains the integrity of the original problem definition. The learned initial guess solely influences the starting point of the search, without altering the underlying problem setup, while MPC acts as a shield for the learned policy so that the control solution satisfies the constraints of the system, which ensures safer operation of the control system. Moreover, the proposed algorithm enables seamless adaptation to new conditions without the need for extensive retraining. During testing, the warm-start policy is tested on both the training domains and some holdout domains, demonstrating the generalizability of our framework in providing a good initial guess. Our key contributions are:
@@
    Propose a two-phase learning framework featuring offline training and online fine-tuning to learn a warm-start policy that provides the iterative control optimization algorithm with higher-quality initial guesses, aiding the command of a high-speed vehicle on multiple novel, challenging Formula 1 tracks in real-time where traditional warm start methods struggle.
@@
    Empirically evaluate our proposed two-phase learning framework and show the online fine-tuning phase helps the iterative control optimization algorithm achieve a $3.9\%$ higher performance in optimization time and a $3.6\%$ higher performance in tracking accuracy on challenging holdout tracks.
@@ II-A Model Predictive Control
Unlike traditional control methods, MPC utilizes a predictive model of the system to anticipate future behavior of the system over a defined planning horizon, $H$. The control problem is formulated as an optimization task, where the objective is to minimize a cost function while satisfying constraints on the system’s inputs and outputs. This optimization problem is typically solved iteratively at each time step, generating a sequence of control inputs that steer the system towards a desired state while considering future predictions of its behavior. When MPC observes a new state, $x_{t}$, at time step, $t$, it optimizes a sequence of control inputs, $U$, over the planning horizon by iteratively minimizing a cost function $J$ while respecting system dynamics and constraints. Then, only the first action in the action sequence is applied to the controlled system, which leads the system to the next state. The optimization process is performed again at the next state. The MPC control law is formulated as shown below in Equations (1):
@@
Solvers used to address the above optimization problem can be divided into gradient-based and gradient-free solvers. Gradient-based solvers leverage the gradient of the cost function with respect to the control inputs to iteratively adjust the control inputs towards the optimal solution [lee2011model, schwenzer2021review]. These solvers are effective when the cost function and constraints are smooth and differentiable, while gradient-free solvers do not rely on gradient information for optimization. Instead, gradient-free solvers explore the search space using heuristics, pattern searches, or stochastic techniques to find the optimal solution [lee2011model, schwenzer2021review]. To relax the constraints imposed on the problem formulation and demonstrate the generalizability of our framework, we choose to use the gradient-free solver COBYLA [powell1994direct] in our experiments.
@@ III Method
In this section, we discuss in detail the proposed two-phase training framework. In the first phase, we run the MPC to collect the expert demonstrations, which are represented as state-action pairs. Then, we use behavior cloning to train a warm-start policy to mimic the expert MPC’s solution, as shown in Algorithm 1. The output of warm-start policy is utilized as an initial guess to warm start the MPC and reduce the optimization time. In the second phase, we load the pre-trained trajectory prediction model into an online training framework and fine tune the warm-start policy to address the suboptimality problem caused by behavior cloning and improve the model’s generalizability. The online fine tuning phase is shown in Algorithm 2.
@@
3:$t\leftarrow 0,{\color[rgb]{0,0,0}s\leftarrow s_{0}},\mathcal{D}=\emptyset$
@@ III-A Offline Training
At this phase, we implement an expert MPC to collect a dataset $\mathcal{D}$ containing $N$ state-action pairs first. The expert MPC, $\pi^{MPC}$ aims to control the agent to finish the task without relying on a warm-started initial guess for the control action sequence. At each step, $\pi^{MPC}$ observes a state $s$ and takes an all-zero vector as the initial guess and optimizes that initial guess to output $(u_{t},u_{t+1},\cdots,u_{t+H-1})$ (line 5). Then the state-action pair, $(s,u_{t},u_{t+1},\cdots,u_{t+H-1})$ is stored in the dataset $\mathcal{D}$ (line 6). The first action in the action sequence is applied to the system, which leads the system to the next state based on the environment transition $T$ (line 7-8). During the data collection stage, $\pi^{MPC}$ optimizes the initial guess for enough iterations to make sure it achieves a high control performance by disabling the “early stop” for the expert MPC.
We design our warm-start policy, $\pi_{\theta}^{warm}$, as a multi-layer perceptron with ReLU activation function [glorot2010understanding]. Given the current state of the vehicle, the $\pi_{\theta}^{warm}$ predicts a sequence of actions which then serves as the initial guess of the MPC to warm start the optimization process. In the offline training phase, we utilize behavior cloning to train $\pi_{\theta}^{warm}$ (line 10).
Behavior Cloning (BC) is a simple yet effective algorithm to learn from demonstrations. The demonstration is a set of trajectories: $\mathcal{D}=\{\tau_{i}\}$. BC learns a control policy, $\pi_{\theta}$, by minimizing the Mean Squared Error (MSE) in the demonstration set, shown in Equation (2). Our algorithm utilizes pre-collected MPC data to perform BC as a warm start for the MPC to improves the MPC’s optimization time.
@@ III-B Online Fine Tuning
Our second phase, online fine tuning, further improves the warm-start policy by learning from the data gathered online using the policy trained through behavior cloning. This phase combines the best aspects of Reinforcement Learning and Dataset Aggregation (DAgger) [ross2011reduction] algorithm. Reinforcement Learning (RL) operates under the formalization of Markov Decision Process (MDP), $\mathcal{M}=\langle\mathcal{S},\mathcal{A},R,T,\gamma,\rho_{0}\rangle$. $\mathcal{S}$ is the state space and $\mathcal{A}$ denotes the action space. $R$ encodes the reward of a given state. $T$ is a deterministic transition function that decides the next state, $s^{\prime}$, when applying the action, $a\in A$, in state, $s\in S$. $\gamma\in(0,1)$ is the temporal discount factor. $\rho_{0}$ denotes the initial state probability distribution. A policy, $\pi:S\rightarrow[0,1]^{|A|}$, is a mapping from states to actions or to a probability distribution over actions. The objective of RL is to find the policy that optimizes the expected discounted return, $\pi^{*}=\mathbb{E}_{\tau\sim\pi}\left[\sum^{\infty}_{t=0}\gamma^{t}R(s_{t})\right]$.
@@
DAgger builds upon behavior cloning (BC) by incorporating online interaction with the environment and online querying of the expert. Unlike BC, which trains solely on a fixed dataset of expert demonstrations, DAgger actively collects data from interactions with the environment and solicits expert feedback to augment its training. This online improvement process allows DAgger to learn from a more diverse set of experiences, adapt to new situations, and refine the agent’s policy over time, ultimately leading to improved performance in imitation learning tasks.
In our framework, RL is used to address the sub-optimality and covariance shift problem caused by offline behavior cloning. DAgger, on the one hand, is adopted to regulate the trajectory prediction model during reinforcement learning training, ensuring it does not forget the experience learned from behavior cloning. Further more, DAgger improves the performance of imitation learning by expanding the diversity of the expert MPC’s demonstrations. Given that offline reinforcement learning algorithms require lots of data to effectively cover the entire state space, we choose to fine-tune our pre-trained trajectory prediction model using Proximal Policy Optimization (PPO) [schulman2017proximal] as our RL training algorithm. The trajectory prediction model acts as a policy network in the PPO framework. DAgger is integrated as a loss term in the RL training, as shown in Equation (3). $L_{RL}$ is the standard PPO loss which includes three parts: $L_{policy}$, $L_{entropy}$, and $L_{value}$. $L_{imitation}$ is the loss signal from DAgger. It represents MSE between the expert MPC’s solution and the warm-started initial guess output from $\pi_{\theta}^{warm}$. $\lambda$ is the weight coefficient for the RL loss and imitation loss.
$L$ $=\lambda\cdot L_{RL}+(1-\lambda)\cdot L_{imitation}$ (3)
$L_{RL}$ $=\lambda_{1}\cdot L_{policy}+\lambda_{2}\cdot L_{entropy}+\lambda_{3}\cdot L_{value}$
@@
In our driving task, the reward at each step of the RL training is shown in Equation (5). The first term is the negative MPC optimization time, and the second term is the negative of the accumulated Cross Track Error ($xte$) over the planning horizon, $H$. The $xte$ is defined in Equation (4). At each step, $\pi_{\theta}^{warm}$ will output an action sequence of length $H$ based on the current state vector, $s_{t}$. Then, the dynamics model, $M_{dynamics}$, is used to calculate the future positions of the agent. The $xte$ is computed between the reference trajectory and the future positions of the vehicle, $pos_{i}^{car}$. The reward obtained at each step directly influences $L_{RL}$ by contributing to the calculation of the advantage estimate, which measures the discrepancy between the observed reward and the expected value of the state-action pair. This estimate affects both $L_{policy}$, which encourages actions leading to higher rewards, and $L_{value}$, which trains the value function to better estimate cumulative rewards.
$r_{t}=-time_{MPC}-\sum_{i=t+1}^{t+H}xte(pos_{i}^{car},Traj^{ref})$ (5)
This reward design helps to optimize the $\pi_{\theta}^{warm}$ by minimizing the MPC running time and by minimizing the difference between the planned trajectory and the reference trajectory, thereby reducing MPC optimization time. This helps to address the sub-optimality problem caused by behavior cloning.

### v2

# Faster Model Predictive Control via Self-Supervised Initialization Learning

@@ Abstract
Optimization for robot control tasks, spanning various methodologies, includes Model Predictive Control (MPC). However, the complexity of the system, such as non-convex and non-differentiable cost functions and prolonged planning horizons often drastically increases the computation time, limiting MPC’s real-world applicability. Prior works in speeding up the optimization have limitations on optimizing MPC running time directly and generalizing to hold out domains. To overcome this challenge, we develop a novel framework aiming at expediting optimization processes directly. In our framework, we combine offline self-supervised learning and online fine-tuning to improve the control performance and reduce optimization time. We demonstrate the success of our method on a novel and challenging Formula 1 track driving task. Comparing to single-phase training, our approach achieves a $19.4\%$ reduction in optimization time and a $6.3\%$ improvement in tracking accuracy on zero-shot tracks.
@@ I INTRODUCTION
Algorithms that optimize control outputs iteratively like MPC have been widely adopted to control dynamic systems, such as autonomous vehicles [1, 2, 3, 4], aircraft [5, 6], humanoid robots [7], etc. However, MPC in real-world applications is challenging due to computational time bottlenecks [8, 9, 10, 11], particularly with complex systems featuring non-convex and non-differentiable cost functions and prolonged planning horizons. [12, 13]. Addressing this challenge is paramount for enhancing the practical utility of MPC in various domains.
@@
Figure 1: Overview of our proposed algorithm. The first two blocks denote the two-phase training framework. In the first phase, we collect expert MPC demonstrations and train a warm-start policy using behavior cloning to speed up MPC. In the second phase, we fine-tune this policy within an online training framework to enhance its performance and generalizability. During testing, the proposed framework is evaluated on both training tracks and challenging zero-shot tracks, as demonstrated in the third block.
@@
Warm starting in robot control optimization involves initializing the solver with a closer-to-optimal solution, speeding up the iterative optimization process. The intuition behind warm starting is that, at each iteration, the solver refines its solution based on previous iterations. As shown in Figure 2, a closer initial guess reduces the distance the optimization algorithm needs to traverse to converge to the optimal solution. By starting closer to the optimum, the solver can often avoid lengthy exploration of the solution space, leading to faster convergence and reduced computational effort. In this paper, we choose to test our proposed algorithm to warm start Model Predictive Control (MPC), which optimizes a trajectory over a predefined horizon.
@@
To provide MPC with a high quality initial guess, various approaches have been adopted. Traditionally, one common technique involves utilizing the MPC solution from the previous sampling instance as the initial guess for the current control step [14, 15]. However, this approach falls short when faced with sudden state changes (e.g., the vehicle approaches a sharp turn). Another method involves maintaining a memory buffer to store historical MPC solutions, from which a suitable initial guess can be retrieved for future planning steps [10, 16]. However, this approach may lack zero-shot generalizability. As the system’s state becomes heterogeneous, the size of the memory buffer needs to increase, thus causing time-consuming operations to searching for the initial guess within the buffer. Alternatively Klaučo et al. propose using a k-NN classifier to classify the solution space into active sets for the solver to search for a solution [17]. However, this method lacks precise warm-start guesses and struggles with complex control tasks involving long planning horizons and heterogeneous observations. Moreover, it’s limited to strictly convex quadratic programs, reducing its applicability to real-world problems with non-linear or non-differentiable objectives and dynamics.
More recent approaches have adopted supervised learning to boost the performance of model predictive optimization algorithms. [18, 19] adopts behavior cloning (BC) to learn the distribution of the initial guesses for the solver. However, supervised learning cannot directly learn to minimize the computation time of MPC. Instead, all supervised learning can do is to imitate the output of the solver. Besides, BC also suffers from covariance shift, which degrades model’s performance on zero-shot scenarios [20]. [21] used reinforcement learning (RL) to train a warm-start policy to help quadrotor track simple zig-zag trajectories. Yet, RL algorithms are too data hungry to learn an optimal policy from scratch, posing challenge to complex, long-horizon tasks [22]. Furthermore, [21] is still not directly learning to minimize the computation time of MPC. Thus, there is a pressing need for innovative approaches that can effectively address the challenges posed by the inherent complexity and non-linearity of real-world control systems.
@@
In this paper, we propose a two-phase learning framework to learn a warm-start policy that provides a better initial guess for iterative control optimization algorithms (such as MPC) to reduce the optimization time. Our proposed algorithm is depicted in Figure 1, which includes a two-phase training framework and an additional zero-shot testing phase. The two training phase works reciprocally to each other. On the one hand, the online-fine tuning phase uses RL to address the suboptimality and covariance shift problem of the offline BC by directly optimize the computation time of MPC. On the other hand, the offline BC provides online learning algorithm with a good starting point to expedite the training process. Moreover, another advantage of using a learned policy to initialize the MPC, rather than for end-to-end control, is that it preserves the original definition of the control problem. The learned policy only affects the starting point, while MPC ensures the control solution satisfies system constraints, leading to safer operation. Our key contributions are:
@@
    Propose a two-phase learning framework featuring offline training and online fine-tuning to learn a warm-start policy that provides the iterative control optimization algorithm with higher-quality initial guesses, aiding the command of a high-speed vehicle on multiple novel, challenging Formula 1 tracks in real-time where traditional warm start methods all failed.
@@
    Empirically evaluate our proposed two-phase learning framework and show the online fine-tuning phase helps the iterative control optimization algorithm achieve a $19.4\%$ higher performance in optimization time and a $6.3\%$ higher performance in tracking accuracy on challenging zero-shot tracks comparing to single-phase training.
@@ II-A Model Predictive Control
MPC is formulated as an optimization task, where the objective is to minimize a cost function while satisfying constraints on the system’s inputs and outputs over a finite horizon, H. This optimization problem is typically solved iteratively at each time step, generating a sequence of control inputs that steer the system towards a desired state while considering future predictions of its behavior. When MPC observes a new state, $x_{t}$, at time step, $t$, it optimizes a sequence of control inputs, $U$, over the planning horizon by iteratively minimizing a cost function , $J$ while respecting system dynamics and constraints. Then, only the first action in the action sequence is applied to the controlled system, which leads the system to the next state. The optimization process is performed again at the next state. The MPC control law is formulated as shown below in Equations (1):
@@ III Method
In this section, we discuss in detail the proposed two-phase training framework. In the first phase, we run the MPC to collect the expert demonstrations, which are represented as state-action pairs. Then, we use BC to train a warm-start policy to mimic the expert MPC’s solution, as shown in Algorithm 1. The output of warm-start policy is utilized as an initial guess to warm start the MPC and reduce the optimization time. In the second phase, we load the pre-trained trajectory prediction model into an online training framework and fine tune the warm-start policy to address the suboptimality problem caused by BC and improve the model’s generalizability. The online fine tuning phase is shown in Algorithm 2.
@@
3:$t\leftarrow 0,s\leftarrow s_{0},\mathcal{D}=\emptyset$
@@ III-A Offline Training
At this phase, we implement an expert MPC to collect a dataset $\mathcal{D}$ containing $N$ state-action pairs first. The expert MPC, $\pi^{MPC}$ aims to control the agent to finish the task without relying on a warm-started initial guess for the control action sequence. At each step, $\pi^{MPC}$ observes a state $s$ and takes an all-zero vector as the initial guess and optimizes that initial guess to output $(u_{t},u_{t+1},\cdots,u_{t+H-1})$ (line 5). Then the state-action pair, $(s,u_{t},u_{t+1},\cdots,u_{t+H-1})$ is stored in the dataset $\mathcal{D}$ (line 6). The first action in the action sequence is applied to the system, which leads the system to the next state based on the environment transition $T$ (line 7-8). During the data collection stage, $\pi^{MPC}$ optimizes the initial guess for enough iterations to ensure it achieves a high control performance by disabling the “early stop” for the expert MPC.
We design our warm-start policy, $\pi_{\theta}^{warm}$, as a multi-layer perceptron with ReLU activation function [23]. Given the current state of the vehicle, the $\pi_{\theta}^{warm}$ predicts a sequence of actions which then serves as the initial guess of the MPC to warm start the optimization process. In the offline training phase, we utilize BC to train $\pi_{\theta}^{warm}$ (line 10).
BC is a simple yet effective algorithm to learn from demonstrations. The demonstration is a set of trajectories: $\mathcal{D}=\{\tau_{i}\}$. BC learns a control policy, $\pi_{\theta}$, by minimizing the Mean Squared Error (MSE) in the demonstration set, shown in Equation (2). Our algorithm utilizes pre-collected MPC data to perform BC as a warm start for the MPC to improves the MPC’s optimization time.
@@ III-B Online Fine Tuning
Our second phase, online fine tuning, further improves the warm-start policy by learning from the data gathered online using the policy trained through BC. This phase combines the best aspects of RL and Dataset Aggregation (DAgger) [24] algorithm. RL operates under the formalization of Markov Decision Process (MDP), $\mathcal{M}=\langle\mathcal{S},\mathcal{A},R,T,\gamma,\rho_{0}\rangle$. $\mathcal{S}$ is the state space and $\mathcal{A}$ denotes the action space. $R$ encodes the reward of a given state. $T$ is a deterministic transition function that decides the next state, $s^{\prime}$, when applying the action, $a\in A$, in state, $s\in S$. $\gamma\in(0,1)$ is the temporal discount factor. $\rho_{0}$ denotes the initial state probability distribution. A policy, $\pi:S\rightarrow A$ , is a mapping from states to actions or to a probability distribution over actions. The objective of RL is to find the policy that optimizes the expected discounted return, $\pi^{*}=\mathbb{E}_{\tau\sim\pi}\left[\sum^{\infty}_{t=0}\gamma^{t}R(s_{t})\right]$.
@@
DAgger builds upon BC by incorporating online interaction with the environment and online querying of the expert. Unlike BC, which trains solely on a fixed dataset of expert demonstrations, DAgger actively collects data from interactions with the environment and solicits expert feedback to augment its training. This online improvement process allows DAgger to learn from a more diverse set of experiences, adapt to new situations, and refine the agent’s policy over time, ultimately leading to improved performance in imitation learning tasks.
In our framework, DAgger and RL address the covariance shift problem caused by offline BC. Additionally, RL not only fine-tunes the warm-start solution but also has the unique advantage of directly optimizing MPC computation time. We choose to leverage Twin Delayed Deep Deterministic Policy Gradient (TD3) [25] as the RL algorithm to fine-tune our pre-trained trajectory prediction model. The trajectory prediction model acts as a policy network in the TD3 framework. DAgger is integrated as a term in actor loss in the RL training, as shown in Equation (3). $B$ is the batch size and $Q_{\theta}$ is the critic network. $L_{imitation}$ is the loss signal from DAgger, representing MSE between the expert MPC’s solution and the warm-started initial guess output from $\pi_{\theta}^{warm}$. Note that DAgger is only added when the quality of the initial guess is lower than a certain threshold. Because we only need to query the expert MPC when the initial guess is of low quality and calling the expert MPC at each step could be time-consuming. $\lambda$ is the weight coefficient between the RL loss and imitation loss.
$L$ $=\lambda\cdot L_{actor}+(1-\lambda)\cdot L_{imitation}$ (3)
$L_{\text{actor}}$ $=-\frac{1}{B}\sum_{i=1}^{B}Q_{\theta}(s_{i},\pi^{\text{warm}}(s_{i}))\cdot L_{value}$
In our driving task, the reward at each step of the RL training is shown in Equation (4). The first term is the negative MPC optimization time, and the second term is the negative of the Cross Track Error ($xte$) over the planning horizon, $H$. The $xte$ is defined in Equation (5). At each step, $\pi_{\theta}^{warm}$ will output an action sequence of length $H$ based on the current state vector, $s_{t}$. Then, the dynamics model, $M_{dynamics}$, is used to calculate the future positions of the agent. The $xte$ is computed between the reference trajectory and the future positions of the vehicle, $pos_{i}^{car}$. The reward obtained at each step directly influences $L_{RL}$ by contributing to the calculation of the advantage estimate, which measures the discrepancy between the observed reward and the expected value of the state-action pair. This estimate affects both $L_{policy}$, which encourages actions leading to higher rewards, and $L_{value}$, which trains the value function to better estimate cumulative rewards.
$r_{t}=-time_{MPC}-xte(pos_{t}^{car},Traj^{ref})$ (4)
@@
This reward design helps to optimize the $\pi_{\theta}^{warm}$ by minimizing the MPC running time and by maximizing the tracking accuracy. This helps to address the sub-optimality problem caused by BC.